OpenStudy (emogirlatemoooocow):

What is the distance between points (9, –7) and (9, 4) on a coordinate plane? units

OpenStudy (solomonzelman):

The distance formula, \(\Large\color{black}{ \displaystyle {\rm D} =\sqrt{ (\color{blue}{{\rm y}_1}-\color{red}{{\rm y}_2})^2+(\color{green}{{\rm x}_1}-\color{darkgoldenrod}{{\rm x}_2})^2 }}\) where \(\Large\color{black}{ \displaystyle {\rm D} }\) is the distance. \(\Large\color{black}{ \displaystyle (\color{green}{{\rm x}_1}~,~~\color{blue}{{\rm y}_1}) }\) and \(\Large\color{black}{ \displaystyle (\color{darkgoldenrod}{{\rm x}_2}~,~~\color{red}{{\rm y}_2}) }\) are your two points.

OpenStudy (anonymous):

That would be 11 units

OpenStudy (anonymous):

Like @SolomonZelman said with that formula you can calculate that it is 11 units

OpenStudy (anonymous):

@EmogirlAtEmoooocow medal? and @ me and fan me for more questions.

OpenStudy (solomonzelman):

I see a policy violation. You can't ask a user to fan yourself like this (explicitly). If the user wants to fan you, he will do so, but you shall not ask! Also, giving out direct answers is not only a policy violation but you are stealing from this questioner - because it is the knowledge that (s)he should acquire.

OpenStudy (solomonzelman):

please, refrain from doing what you did in the future. Unfamiliar with the policy? Then read \(\bf\color{re}{\href{http:///openstudy.com/code-of-conduct}{Code~of~Conduct}}\) .

OpenStudy (emogirlatemoooocow):

thanks @SolomonZelman

OpenStudy (solomonzelman):

sure, yw